# Higher-dimensional Teter rings via the canonical trace

### v1

@@ Abstract.
We study Puthenpurakal’s higher-dimensional Teter rings via the canonical trace ideal. We give a sufficient criterion for Teterness and show that, in the standard graded case, it is also necessary, yielding a characterization. Consequently, several nearly Gorenstein families are Teter; moreover, under certain hypotheses, the Cohen–Macaulay type of nearly Gorenstein rings is bounded by the codimension. We also analyze Teterness for fiber products, Veronese subrings, and numerical semigroup rings.
@@ 1. Introduction
We say $R$ is Teter if there exists a graded RR–homomorphism $\varphi:\ \omega_{R}\rightarrow R$ such that either $\varphi(\omega_{R})={\mathfrak{m}}$ (this happens only when $\dim(R)=0$), or $\varphi$ is injective and $\operatorname{embdim}\!\bigl(R/\varphi(\omega_{R})\bigr)\leq\dim R$.
It recovers the classical Artinian case at $\dim R=0$ and, in positive dimension, aligns with [puthenpurakal2025higher]: Teter in our sense implies $R_{{\mathfrak{m}}}$ is Teter in theirs. If $R$ is a domain, the notions are equivalent, i.e., $R$ is Teter if and only if $R_{{\mathfrak{m}}}$ is Teter. Thus Section  1 furnishes a natural graded analogue.
The canonical trace $\operatorname{tr}_{R}(\omega_{R})$ is the ideal generated by the images of all maps $\omega_{R}$ to $R$; it detects the non–Gorenstein locus and has seen renewed interest. We establish sufficient conditions for $R$ to be Teter, phrased in terms of the canonical trace; moreover, in the standard graded case these conditions characterize Teter rings. We fix minimal notation: Set $r_{0}(R):=\dim_{R_{0}}([\omega_{R}]_{-a_{R}})$. $R$ is called level ([stanley2007combinatorics]) if $r(R)=r_{0}(R)$, where $r(R)$ is the Cohen–Macaulay type and $a_{R}$ the $a$-invariant of $R$. Let $\text{indeg}({\mathfrak{m}}):=\operatorname{min}\{i\in\mathbb{Z}:[\mathfrak{m}_{R}]_{i}\neq 0\}$ and let $\text{codim}(R):=\operatorname{embdim}(R)-\dim(R)$. Our main result is stated as follows.
@@
Assume that $\dim(R)>0$. Consider the following conditions:
@@
    $\omega_{R}]_{-a_{R}}$ contains a torsion-free element $x$ of $\omega_{R}$ (that is, $rx=0$ implies $r=0$ for all $r\in R$), $r_{0}(R)\geq\text{codim}(R)$ and $\operatorname{tr}(\omega_{R})]_{\text{indeg}({\mathfrak{m}})}$ contains a non-zerodivisor;
@@
    $R$ is a level Teter ring with $r(R)=\text{codim}(R)$;
@@
    $R$ is Teter ((in which case $r(R)=\operatorname{codim}(R)$)).
Then $(1)\Rightarrow(2)\Rightarrow(3)\Rightarrow(4)$ holds. Moreover, if $R$ is standard graded, all the conditions are equivalent.
We present some applications of Theorem 1.2. It is known that a Cohen–Macaulay graded ring is Gorenstein if and only if the trace ideal of its canonical module coincides with the whole ring (see [herzog2019trace, Lemma 2.1]). As a generalization of the Gorenstein property based on this fact, [herzog2019trace] introduced the notion of a nearly Gorenstein Cohen–Macaulay ring, namely a Cohen–Macaulay ring whose trace ideal of the canonical module contains the (graded) maximal ideal.
Gorenstein rings are characterized as Cohen–Macaulay rings of Cohen–Macaulay type one. Concerning the Cohen–Macaulay type of nearly Gorenstein rings, several results are known, in particular when the codimension is small. For instance, if the codimension is $2$ and the Krull dimension is positive, then the Cohen–Macaulay type is at most $2$. This follows immediately as a consequence of an explicit formula for the canonical trace obtained by applying the Hilbert–Burch theorem, the structure theorem for minimal free resolutions (see [ficarra2024canonical!, Theorem 1.1] or [ficarra2025canonical, Theorem 2.1]).
When the codimension is at least $3$, no tool as powerful as the Hilbert–Burch theorem is available, and hence computing and analyzing the canonical trace from a minimal free resolution becomes much more difficult than in the codimension $2$ case. On the other hand, in the case of nearly Gorenstein numerical semigroup rings of codimension $3$ or $4$, various techniques have been developed, and several interesting results on the Cohen–Macaulay type are known. For instance, [moscariello2021nearly, Theorem 2.4] proves that a nearly Gorenstein monomial curve of codimension three has Cohen–Macaulay type at most $3$. In more general situations with embedding dimension at least $4$, Stamate posed the following question ([stamate2016betti, Question 3.7]): “Is there any bound for the type of the numerical semigroup ring in terms of the embedding dimension of $R$ when $R$ is nearly Gorenstein?”
On the other hand, for several classes of rings other than numerical semigroup rings, it is known that if the codimension is $3$ and the ring is nearly Gorenstein, then the Cohen–Macaulay type is at most $3$. For example, if a simplicial affine semigroup ring is nearly Gorenstein of codimension $3$, then its Cohen–Macaulay type is at most $3$ ([jafari2024nearly, Theorem 3.9]). In particular, for (non-Gorenstein) projective monomial curves which are $2$-dimensional and standard graded, it is known that the Cohen–Macaulay type is exactly $3$ ([miyashita2023nearly, Theorem A]). Furthermore, for non-Gorenstein nearly Gorenstein Stanley–Reisner rings of codimension $n$, it also follows immediately from [miyashita2025canonical, Theorem A] that the type is equal to $n$. In view of the above, we are led to the following natural problem for higher-dimensional rings.
@@
If $R$ is nearly Gorenstein and ${\text{codim}}(R)=3$, then is it true that $r(R)\leq 3$? Moreover, is there a bound for the type of $R$ in terms of the codimension of $R$ when $R$ is nearly Gorenstein?
By applying Theorem 1.2, we give an affirmative answer to Section  1 under suitable assumptions. Recall that $r_{0}(R):=\dim_{R_{0}}([\omega_{R}]_{-a_{R}})$.
@@
    If $r_{0}(R)\geq\text{codim}(R)$, then $R$ is level and $r(R)=\text{codim}(R)$;
@@
    If $R$ is standard graded, $\text{codim}(R)=3$, $\dim(R)\geq 2$ and $r_{0}(R)\neq 2$, then $R$ is level and $r(R)=3$;
Some of the conclusions of Section  1 in fact hold without assuming that $R$ is nearly Gorenstein; they remain valid under weaker hypotheses on the canonical trace (see Section  4).
In [puthenpurakal2025higher], it is proved that several classes of rings are Teter. For instance, the completion of any one-dimensional standard graded generically Gorenstein ring with minimal multiplicity over an infinite field, as well as the completion of suitable graded rings of finite Cohen–Macaulay type, is shown to be Teter (see [puthenpurakal2025higher, Theorem 1.4, Proposition 5.2]). Combining Theorem 1.2 with known facts about the canonical trace, we show that a certain known class of (non-Gorenstein) nearly Gorenstein standard graded rings is Teter. Moreover, tensor products, fiber products, and Segre products are basic operations in ring theory, and we also show that they can be used to construct new examples of Teter rings.
Corollary 1.5 (A part of Section  4 and § 5.1 and § 5.2).
@@
    Assume that $R$ is a standard graded nearly Gorenstein ring. Then it is Teter, if it satisfies one of the following conditions:
@@
      $R$ has minimal multiplicity (i.e. $e(R)=\operatorname{embdim}(R)-\dim(R)+1$);
@@
      $R$ is a Stanley–Reisner ring.
@@
    Assume that $A$ and $B$ are Cohen–Macaulay generically Gorenstein standard graded rings with $\dim(A)=\dim(B)=1$. Furthermore, suppose that at least one of $A$ and $B$ is not regular. Then the fiber product $A\times_{\mathbb{k}}B$ of $A$ and $B$ over a field $\mathbb{k}$ (see § 5.1) is Teter if and only if both $A$ and $B$ have minimal multiplicity;
@@
    Assume that $R$ is a (non-Gorenstein) generically Gorenstein standard graded ring with minimal multiplicity. If either $\dim(R)=1$ or $\operatorname{tr}(\omega_{R})={\mathfrak{m}}$ with $\dim(R)=2$, then the $k$-th Veronese subring of $R$ is Teter for every $k\geq 2$.
Finally, we investigate the Teter property of numerical semigroup rings $R_{H}$, which are a typical class of rings that are not even semi-standard graded. We show that the Teter property of numerical semigroup rings can be characterized in terms of the minimal system of generators and the pseudo-Frobenius numbers of the numerical semigroup (see Theorem 6.5). Using this characterization, we obtain the following results for numerical semigroups with minimal multiplicity or codimension $2$.
Theorem 1.6 (Section  6, Section  6 and Section  6).
Let $H$ be a numerical semigroup and let $R_{H}$ denote its numerical semigroup ring. Assume $R_{H}$ is not Gorenstein. Then the following hold:
  1. (a)
    Suppose $R_{H}$ has minimal multiplicity. If either ${\text{codim}}(R_{H})=2$ or $\operatorname{tr}(\omega_{R_{H}})]_{a_{1}}\neq 0$, then $R_{H}$ is Teter.
  2. (b)
    Suppose ${\text{codim}}(R_{H})=2$. If $R_{H}$ is nearly Gorenstein but not almost Gorenstein, then it is Teter. If $R_{H}$ is almost Gorenstein, the Hilbert–Burch matrix of $R_{H}$ can be expressed as $A=\begin{pmatrix}x_{1}^{a}&x_{2}^{b}&x_{3}^{c}\\ x_{2}&x_{3}&x_{1}\end{pmatrix}$ where $a,b,c$ is positive integers. In this case, $R_{H}$ is Teter if and only if $\operatorname{min}\{a,b,c\}=1$.
@@
In Section  2, we review basic definitions and properties of Cohen–Macaulay graded rings, trace ideals, and numerical semigroup rings. Section  3 is devoted to the Teter property of graded rings: we prove our main result Theorem 3.13 and study Teter rings in the (semi-standard) graded setting. In Section  4 and Section  5, we relate nearly Gorenstein and Teter rings and study the behavior of the Teter property under constructions such as fiber products and Veronese subrings. Finally, in Section  6, we characterize the Teter property of numerical semigroup rings (see Theorem 6.5), obtaining results for graded rings that are not even semi-standard graded.
Setup 1.7.
@@ 2.1. Canonical modules over Cohen–Macaulay graded rings
In this paper, for simplicity of notation, we denote the Cohen–Macaulay type of $R$ by $r(R)$, and set $r_{0}(R):=\dim_{R_{0}}([\omega_{R}]_{-a_{R}})$. Note that $r_{0}(R)\leq r(R)$ always holds. $R$ is called level ([stanley2007combinatorics]) if $\omega_{R}$ is generated in a single degree; equivalently, $r(R)=r_{0}(R)$ (see, e.g., [stanley2007combinatorics, Proposition 3.2]). Moreover, following the terminology of [ene2015pseudo], we say that $R$ is pseudo-Gorenstein if $r_{0}(R)=1$.
@@ 2.2. Semi-standard graded rings, Hilbert series, and almost Gorenstein graded rings
${\text{Hilb}}(R,t)=\frac{h_{0}+h_{1}t+\cdots+h_{s}t^{s}}{(1-t)^{\dim R}},$
@@
Assume that $R$ is semi-standard graded, and let $h(R)=(h_{0},h_{1},\ldots,h_{s(R)})$ be its $h$-vector. We always have $h_{0}=1$ since $h_{0}={\text{Hilb}}(R,0)=\dim_{R_{0}}(R_{0})=1$. Moreover, if $R$ is Cohen–Macaulay, its $h$-vector satisfies $h_{s(R)}=r_{0}(R)$ and $h_{i}\geqq 0$ for every $0\leq i\leq s(R)$. For further background on $h$-vectors of Noetherian graded rings, see [stanley1978hilbert, stanley2007combinatorics].
@@ 2.3. Trace ideals over graded rings
Remark 2.8 ([kumashiro2025canonical, Remark 2.2]).
Since $R$ is Noetherian, we have
$\operatorname{tr}_{R}(M)=\sum_{\varphi\in\operatorname{Hom}_{R}(M,R)}\varphi(M).$
@@
The next result is a generalization of [miyashita2024canonical, Proposition 3.7]. Its proof is essentially the same as that of [miyashita2024canonical, Proposition 3.7], but we include it for the reader’s convenience.
@@
Let $R$ be a Noetherian graded ring and let $J\subset{}^{*}Q(R)$ be a fractional ideal of $R$. Suppose that $J_{\text{indeg}_{R}(J)}$ contains an $R$-torsion free element of $J$. If $\operatorname{tr}(J)]_{\text{indeg}_{R}({\mathfrak{m}})}$ contains a nonzero-divisor, then there exists a graded ideal $I\subset R$ such that $J\cong I$ and $\dim_{R_{0}}(I_{\text{indeg}_{R}({\mathfrak{m}})})=\dim_{R_{0}}\left(J_{\text{indeg}_{R}(J)}\right)$.
@@
We may assume that $R_{0}$ is infinite. Assume that there exists a non-zerodivisor $z\in[\operatorname{tr}(J)]_{\text{indeg}_{R}({\mathfrak{m}})}$. Put $n=\dim_{R_{0}}(J_{\text{indeg}_{R}(J)})$ and $a=\text{indeg}_{R}(J)$. Then by [miyashita2024linear, Lemma 2.12 (2)], we may assume that $J$ is minimally generated by $R$-regular elements $f_{1},\cdots,f_{m}$, where $m\geq n$. We may assume that $f_{1},\cdots,f_{n}\in J_{a}$. Note that $\operatorname{tr}(J)=J\cdot J^{-1}$ by [miyashita2024linear, Lemma 2.6]. Since $z\in[\operatorname{tr}(J)]_{\text{indeg}_{R}({\mathfrak{m}})}$, we can write $z=\sum_{i=1}^{m}{f_{i}g_{i}}$, where $g_{1},\cdots,g_{m}$ are homogeneous elements of $J^{-1}$. Moreover, we have $g_{i}=0$ for any $n<i\leq m$ because $f_{1},\cdots,f_{n}\in J_{a}$ are $R$-regular. Therefore the vector space $V=[J^{-1}]_{\text{indeg}_{R}({\mathfrak{m}})-a}$ is equal to $\frac{1}{b}\langle x_{1},\cdots,x_{s}\rangle$ where $b$ is a non-zerodivisor of $R$, $s$ is natural number and $x_{1},\cdots,x_{s}$ homogeneous elements of $R$ of the same degree, $\text{indeg}_{R}({\mathfrak{m}})-a+\operatorname{deg}(b)$.

### v2

@@ Abstract.
We study Puthenpurakal’s higher-dimensional Teter rings via the canonical trace ideal. We give a sufficient criterion for Teterness and show that, in the standard graded case, it is also necessary, yielding a characterization. Consequently, several nearly Gorenstein families are Teter; moreover, under certain hypotheses, the Cohen–Macaulay type of nearly Gorenstein rings is bounded by the codimension.
@@ 1. Introduction
We say $R$ is Teter if there exists a graded RR–homomorphism $\varphi:\ \omega_{R}\rightarrow R$ such that either $\varphi(\omega_{R})={\mathfrak{m}}$ (this happens only when $\dim(R)=0$), or $\varphi$ is injective and $\operatorname{embdim}\!\bigl(R/\varphi(\omega_{R})\bigr)\leq\dim(R)$.
It recovers the classical Artinian case at $\dim(R)=0$ and, in positive dimension, aligns with [puthenpurakal2025higher]: Teter in our sense implies $R_{{\mathfrak{m}}}$ is Teter in theirs. If $R$ is a domain, the notions are equivalent, i.e., $R$ is Teter if and only if $R_{{\mathfrak{m}}}$ is Teter. Thus Section  1 furnishes a natural graded analogue.
The canonical trace $\operatorname{tr}_{R}(\omega_{R})$ is the ideal generated by the images of all maps $\omega_{R}$ to $R$; it detects the non–Gorenstein locus and has seen renewed interest. We establish sufficient conditions for $R$ to be Teter, phrased in terms of the canonical trace; moreover, in the standard graded case these conditions characterize Teter rings. We fix minimal notation: Set $r_{\operatorname{min}}(R):=\dim_{R_{0}}([\omega_{R}]_{-a_{R}})$. $R$ is called level ([stanley2007combinatorics]) if $r(R)=r_{\operatorname{min}}(R)$, where $r(R)$ is the (Cohen–Macaulay) type and $a_{R}$ the $a$-invariant of $R$. Let $\text{indeg}({\mathfrak{m}}):=\operatorname{min}\{i\in\mathbb{Z}:[\mathfrak{m}_{R}]_{i}\neq 0\}$ and let $\text{codim}(R):=\operatorname{embdim}(R)-\dim(R)$. Our main result is stated as follows.
@@
Consider the following conditions:
@@
    $\omega_{R}]_{-a_{R}}$ contains a torsion-free element $x$ of $\omega_{R}$ (that is, $rx=0$ implies $r=0$ for all $r\in R$), $r_{\operatorname{min}}(R)\geq\text{codim}(R)$ and $\operatorname{tr}(\omega_{R})]_{\text{indeg}({\mathfrak{m}})}$ contains a non-zerodivisor;
@@
    $R$ is a Teter and level ring with $r(R)=\text{codim}(R)$ and $\dim(R)>0$;
@@
    $R$ is Teter with $\dim(R)>0$.
Then $(1)\Rightarrow(2)\Rightarrow(3)\Rightarrow(4)$ holds. Moreover, if $R$ is standard graded, all the conditions are equivalent; in particular, if any one of them holds, then $r(R)=r_{\operatorname{min}}(R)=\text{codim}(R)$.
In this section, we give several applications of Theorem 1.2. In [herzog2019trace], Herzog et al. introduced the notion of a nearly Gorenstein ring by studying the trace ideal of the canonical module. While Gorenstein rings are precisely the Cohen–Macaulay rings of type $1$, the type of a nearly Gorenstein ring is often bounded above by the codimension of $R$.
For instance, if ${\text{codim}}(R)=2$ and $\dim(R)>0$, then [ficarra2025canonical, Theorem 2.1] implies that $r(R)\leq 2$. When ${\text{codim}}(R)=3$, it is known that $r(R)\leq 3$ for nearly Gorenstein simplicial affine semigroup rings ([jafari2024nearly]). Moreover, for nearly Gorenstein projective monomial curves one has $r(R)=3$ ([miyashita2023nearly]). In addition, nearly Gorenstein monomial curves in codimension $3$ satisfy $r(R)\leq 3$ ([moscariello2021nearly]).
In [stamate2016betti, Question 3.7], Stamate asked whether the type of a nearly Gorenstein numerical semigroup ring admits an upper bound depending only on the embedding dimension (equivalently, on the codimension). Motivated by the above results, we pose the following natural question in a broader setting:
@@
Let $R$ be a nearly Gorenstein ring. Is its type $r(R)$ bounded above by ${\text{codim}}(R)$? In particular, if $R$ is nearly Gorenstein and ${\text{codim}}(R)=3$, does it follow that $r(R)\leq 3$?
By applying Theorem 1.2 together with several known results, we answer Section  1 affirmatively under suitable hypotheses. Recall the notation $r_{\operatorname{min}}(R):=\dim_{R_{0}}([\omega_{R}]_{-a_{R}})$.
@@
    If $r_{\operatorname{min}}(R)\geq\text{codim}(R)$, then $R$ is level and $r(R)=\text{codim}(R)$;
@@
    If $R$ is standard graded, $\text{codim}(R)=3$, $\dim(R)\geq 2$ and $r_{\operatorname{min}}(R)\neq 2$, then $R$ is level and $r(R)=3$;
In [puthenpurakal2025higher, Section 5], it is shown that several classes of rings, including rings of finite Cohen–Macaulay type and one-dimensional rings of minimal multiplicity, are Teter in the higher-dimensional sense. As an application of Theorem 1.2, we show that various standard graded nearly Gorenstein rings are Teter. Furthermore, we show that the Teter property can be constructed and is preserved under operations such as fiber products and certain Veronese subalgebras.
Corollary 1.5 (See a part of Section  4 and § 5.1 and § 5.2).
@@
    Assume that $R$ is a standard graded nearly Gorenstein. Then $R$ is Teter, if it satisfies one of the following conditions:
@@
      $R$ has minimal multiplicity;
@@
      $R$ is a Stanley–Reisner ring;
@@
    Let $A$ and $B$ be 1-dimensional Cohen–Macaulay generically Gorenstein standard graded rings, at least one of which is not regular. Then the fiber product $A\times_{\mathbb{k}}B$ (see § 5.1)is Teter if and only if both $A$ and $B$ have minimal multiplicity.
@@
    Assume $R$ is a (non-Gorenstein) generically Gorenstein standard graded ring with minimal multiplicity. If either $\dim(R)=1$ or $\operatorname{tr}(\omega_{R})={\mathfrak{m}}$ with $\dim(R)=2$, then $R^{(k)}$ is Teter for every $k\geq 2$.
While standard graded Teter rings satisfy $r(R)=r_{\operatorname{min}}(R)$ by Theorem 1.2, this equality fails for semi-standard graded rings. Specifically, for any $a\in\mathbb{Z}_{>0}$, there exists a nearly Gorenstein Teter ring $R$ such that $r(R)-r_{\operatorname{min}}(R)=a$ (Section  4).
Finally, we investigate the Teter property of numerical semigroup rings, which are a typical class of rings that are not even semi-standard graded. As a main result of this section, we show that the Teter property of numerical semigroup rings can be characterized by using the minimal set of generators and invariants determined by the numerical semigroup called pseudo-Frobenius numbers (see Theorem 6.5). Using this characterization, we obtain the following assertions for the cases where the numerical semigroup has minimal multiplicity or the codimension is $2$ (see Section  6 and Section  6). Furthermore, we obtain the following result analogous to Section  4 (see Section  6).
@@
This paper is organized as follows. Section  2 reviews the basics of Cohen–Macaulay graded rings, trace ideals, and numerical semigroup rings. In Section  3, we present our main result on the Teter property for (semi-)standard graded rings. Section  4 explores the relationship between nearly Gorenstein and Teter rings, establishing sufficient conditions. Section  5 examines how the Teter property behaves under ring operations, specifically fiber products and Veronese subalgebras. Finally, Section  6 provides a characterization of the Teter property for numerical semigroup rings.
Setup 1.6.
@@ 2.1. Canonical modules over Cohen–Macaulay graded rings
In this paper, for simplicity of notation, we denote the (Cohen–Macaulay) type of $R$ by $r(R)$, and set $r_{\operatorname{min}}(R):=\dim_{R_{0}}([\omega_{R}]_{-a_{R}})$. Note that $r_{\operatorname{min}}(R)\leq r(R)$ always holds. $R$ is called level ([stanley2007combinatorics]) if $\omega_{R}$ is generated in a single degree; equivalently, $r(R)=r_{\operatorname{min}}(R)$ (see, e.g., [stanley2007combinatorics, Proposition 3.2]). Moreover, following the terminology of [ene2015pseudo], we say that $R$ is pseudo-Gorenstein if $r_{\operatorname{min}}(R)=1$.
@@ 2.2. Semi-standard graded rings, Hilbert series, and almost Gorenstein graded rings
${\text{Hilb}}(R,t)=\frac{h_{0}+h_{1}t+\cdots+h_{s}t^{s}}{(1-t)^{\dim(R)}},$
@@
Assume that $R$ is semi-standard graded, and let $h(R)=(h_{0},h_{1},\ldots,h_{s(R)})$ be its $h$-vector. We always have $h_{0}=1$ since $h_{0}={\text{Hilb}}(R,0)=\dim_{R_{0}}(R_{0})=1$. Moreover, if $R$ is Cohen–Macaulay, its $h$-vector satisfies $h_{s(R)}=r_{\operatorname{min}}(R)$ and $h_{i}\geq 0$ for every $0\leq i\leq s(R)$. For further background on $h$-vectors of Noetherian graded rings, see [stanley1978hilbert, stanley2007combinatorics].
@@ 2.3. Trace ideals over graded rings
The next result is a generalization of [miyashita2025canonical, Proposition 3.7]. Its proof is essentially the same as that of [miyashita2025canonical, Proposition 3.7], but we include it for the reader’s convenience.
@@
Let $R$ be a Noetherian graded ring and let $J\subset{}^{*}Q(R)$ be a fractional ideal of $R$. Suppose that $J_{\text{indeg}_{R}(J)}$ contains an $R$-torsion free element of $J$. If $\operatorname{tr}(J)]_{\text{indeg}_{R}({\mathfrak{m}})}$ contains a non-zero divisor, then there exists a graded ideal $I\subset R$ such that $J\cong I$ and $\dim_{R_{0}}(I_{\text{indeg}_{R}({\mathfrak{m}})})=\dim_{R_{0}}\left(J_{\text{indeg}_{R}(J)}\right)$.
@@
We may assume that $R_{0}$ is infinite. Assume that there exists a non-zero divisor $z\in[\operatorname{tr}(J)]_{\text{indeg}_{R}({\mathfrak{m}})}$. Put $n=\dim_{\mathbb{k}}(J_{\text{indeg}_{R}(J)})$ and $a=\text{indeg}_{R}(J)$. Then by [miyashita2024linear, Lemma 2.12 (2)], we may assume that $J$ is minimally generated by $R$-regular elements $f_{1},\cdots,f_{m}$, where $m\geq n$. We may assume that $f_{1},\cdots,f_{n}\in J_{a}$. Note that $\operatorname{tr}(J)=J\cdot J^{-1}$ by [miyashita2024linear, Lemma 2.6]. Since $z\in[\operatorname{tr}(J)]_{\text{indeg}_{R}({\mathfrak{m}})}$, we can write $z=\sum_{i=1}^{m}{f_{i}g_{i}}$, where $g_{1},\cdots,g_{m}$ are homogeneous elements of $J^{-1}$. Moreover, we have $g_{i}=0$ for any $n<i\leq m$ because $f_{1},\cdots,f_{n}\in J_{a}$ are $R$-regular. Therefore the vector space $V=[J^{-1}]_{\text{indeg}_{R}({\mathfrak{m}})-a}$ is equal to $\frac{1}{b}\langle x_{1},\cdots,x_{s}\rangle$ where $b$ is a non-zero divisor of $R$, $s$ is natural number and $x_{1},\cdots,x_{s}$ homogeneous elements of $R$ of the same degree, $\text{indeg}_{R}({\mathfrak{m}})-a+\operatorname{deg}(b)$.